# Hybrid Beamforming Design for RSMA-assisted mmWave Integrated Sensing and Communications

### v1

@@ III-C Proposed Iterative Algorithm for Solving Problem P2-3
Algorithm 1 summarizes the proposed CCCP-based iterative algorithm for solving the inner loop problem P2-3, where we implement the above five updating steps in each iteration.
Algorithm 1 CCCP-Based Algorithm in the Inner Loop
1:  Define the accuracy tolerance $\epsilon^{in}$ and the maximum number of iterations $N_{max}^{in}$. Initialize the iteraion number $t=0$.
@@
3:     Update $\{\boldsymbol{\rm w}^{(t)}\}$ according to step 1;
4:     Update $\{\boldsymbol{\rm u}^{(t)}\}$ according to step 2;
5:     Update $\{\boldsymbol{\rm Y}^{(t)},\boldsymbol{\rm Z}_{k}^{(t)},\boldsymbol{\rm q}_{k}^{(t)},\boldsymbol{\rm F}_{\rm D}^{(t)}\}$ according to step 3;
6:     Update $\{\boldsymbol{\rm c}^{(t)},\boldsymbol{\rm F}_{\rm RF}^{(t)}\}$ according to step 4;
7:     Update $\{\boldsymbol{\rm X}^{(t)}\}$ according to step 5;
8:     $t\leftarrow t+1$;
9:  until The difference between the OF values of two adjacent iterations is less than $\epsilon^{in}$ or $t\geq N_{max}^{in}$.
III-D Summary of the Proposed WMMSE-PDD Algorithm
After each inner iteration of Algorithm 1 is completed, we need to update the Lagrange multipliers and the penalty parameter of the outer loop so that the PDD-based algorithm can converge stably. We first define the constraint violation in the $m$-th outer iteration, denoted by $h\left(\mathbb{X}^{(m)}\right)$, expressed as follows:
@@ III-D Summary of the Proposed WMMSE-PDD Algorithm
$\|\boldsymbol{\rm X}-\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm F}_{\rm D}\|_{F}\},k\in\mathcal{K}.$ (51)
In the $(m+1)$-th outer iteration, we introduce the variable $\zeta^{(m+1)}$ associated with the constraint violation $h(\mathbb{X}^{(m)})$, usually set to $\zeta^{(m+1)}=0.9h(\mathbb{X}^{(m)})$, and then give the update rule as follows: when $h(\mathbb{X}^{(m+1)})\geq\zeta^{(m+1)}$, the penalty parameter is updated according to $\rho^{(m+1)}=c_{\rho}\rho^{(m)}(0<c_{\rho}<1)$; otherwise, the Lagrange multipliers need to be updated, which can be given by
$\boldsymbol{\rm\Lambda}_{1}^{(m+1)}=\boldsymbol{\rm\Lambda}_{1}^{(m)}+\frac{\boldsymbol{\rm X}^{(m)}-\boldsymbol{\rm Y}^{(m)}}{\rho^{(m)}},$ (52a)
$\boldsymbol{\rm\Lambda}_{2,k}^{(m+1)}=\boldsymbol{\rm\Lambda}_{2,k}^{(m)}+\frac{\boldsymbol{\rm X}^{(m)}-\boldsymbol{\rm Z}_{k}^{(m)}}{\rho^{(m)}},k\in\mathcal{K},$ (52b)
$\boldsymbol{\rm\Lambda}_{3}^{(m+1)}=\boldsymbol{\rm\Lambda}_{3}^{(m)}+\frac{\boldsymbol{\rm X}^{(m)}-\boldsymbol{\rm F}_{\rm RF}^{(m)}\boldsymbol{\rm F}_{\rm D}^{(m)}}{\rho^{(m)}},$ (52c)
$\boldsymbol{\rm\lambda}_{c,k}^{(m+1)}=\boldsymbol{\rm\lambda}_{c,k}^{(m)}+\frac{\boldsymbol{\rm c}^{(m)}-\boldsymbol{\rm q}_{k}^{(m)}}{\rho^{(m)}},k\in\mathcal{K}.$ (52d)
In summary, the detailed steps of the proposed WMMSE-PDD algorithm for handling joint common rate allocation and hybrid beamforming design are listed in Algorithm 2.
1:  Define the accuracy tolerance $\epsilon^{out}$ and the maximum number of iterations $N_{max}^{out}$. Initialize $\mathbb{X}^{(0)}$, $\boldsymbol{\rm\Lambda}^{(0)}$, $\rho^{(0)}$, $\zeta^{0}$, set $10<c_{\rho}<$ 1, the iteraion number $m=0$.
2:  repeat
3:     Update $\mathbb{X}^{(m+1)}$ according to Algorithm 1;
4:     if $h(\mathbb{X}^{(m+1)})\leq\zeta^{(m)}$ then
5:        Update $\boldsymbol{\rm\Lambda}^{(m+1)}$ by applying Eqs.(52a)-(52d),$\rho^{(m+1)}=\rho^{(m)}$;
@@
Convergence Analysis: First, we have the proposition for problem P2-2 as follows:
Proposition 5
The Mangasarian-Fromovitz constraint qualification (MFCQ) condition for problem P2-2 is satisfied by any feasible solution.
[…]
IV ISAC system design with finite resolution phase shifters
So far, we have designed algorithms for mmWave ISAC systems with infinite-resolution PSs. However, it is difficult to design ultra-high-resolution PSs with existing semiconductor technologies, especially for massive MIMO systems, high-resolution PSs lead to higher hardware and power costs. Hence, we extend the algorithm to finite-resolution PSs to balance the performance and hardware cost of the system.
Specifically, according to [45], the set of finite resolution PS values is defined as $\mathcal{F}\triangleq\{1,e^{j\Delta},e^{j2\Delta},\ldots,e^{j\left(2^{b}-1\right)\Delta}\}$, where $b$ denotes the number of phase quantization bits and $\Delta\triangleq\frac{2\pi}{2^{b}}$ denotes the quantization step. The problem of joint common rate allocation and HBF design with finite-resolution PSs can be formulated as follows:
[…]
s.t. $\left|\boldsymbol{\rm F}_{\rm RF}(i,j)\right|\in\mathcal{F},\forall i,j,$ (53b)
[…]
Based on the fact that the PDD framework can solve distributed optimization problems, we still deal with problem P3 using the WMMSE-PDD algorithm. It is only necessary to modify the subproblem for the variable $\boldsymbol{\rm F}_{\rm RF}$ in Step 4 of Section III-B as follows:
$\min_{\left|\boldsymbol{\rm F}_{\rm RF}(i,j)\right|\in\mathcal{F},\forall i,j}$ $\operatorname{tr}\left(\boldsymbol{\rm F}_{\rm RF}^{H}\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm C}\right)-2\mathfrak{R}\left\{\operatorname{tr}\left(\boldsymbol{\rm F}_{\rm RF}^{H}\boldsymbol{\rm B}\right)\right\}.$ (54)
We can modify [43, Algorithm 4] to solve problem (54), specifically, only replace Step 4 in [43, Algorithm 4] with the following problem:
$x^{(t+1)}=\arg\max_{\left|\boldsymbol{\rm F}_{\rm RF}(i,j)\right|\in\mathcal{F}}$$\mathfrak{R}\left\{(a^{(t)})^{*}\boldsymbol{\rm F}_{\rm RF}(i,j)\right\},$ (55)
where $a^{(t)}=\boldsymbol{\rm F}^{(t)}_{\rm RF}(i,j)\boldsymbol{\rm C}(j,j)-(\boldsymbol{\rm F}^{(t)}_{\rm RF}\boldsymbol{\rm C})(i,j)+\boldsymbol{\rm B}(i,j)$ is the constant calculated from the result of the $t$-th iteration. Further, problem (55) can be solved by a one-dimensional exhaustive search in the set $\mathcal{F}$. As a result, the complexity of solving problem (55) is $\mathcal{O}\left(I_{\rm RF}N_{t}N_{\rm RF}(N_{t}N_{\rm RF}+2^{b})\right)$. Accordingly, the complexity of the WMMSE-PDD algorithm with finite resolution PSs is $\mathcal{O}\big{(}N_{max}^{out}N_{max}^{in}\big{(}(N_{t}(K+1))^{3}+I_{\rm RF}N_{t}N_{\rm RF}(N_{t}N_{\rm RF}+2^{b})+K\log_{2}(\frac{I_{0}}{\epsilon})\big{)}\big{)}.$
V Simulation Results
In this section, we provide simulation results to verify the performance of the proposed HBF for the RSMA-assisted mmWave ISAC system. We consider an RSMA-assisted mmWave ISAC system at the carrier frequency of 28 $\mathrm{GHz}$, where the BS is equipped with $N_{t}=32$ transmit antennas and $N_{r}=32$ receive antennas while serving $K=4$ downlink single-antenna users and detecting a radar target of interest in the presence of clutter interference. Specifically, the users and the radar target are located within 200 $\mathrm{m}$ of the BS. Set the target of interest located at spatial angle $\theta_{0}=10^{\circ}$ and four clutter interferences located at spatial angles $\theta_{1}=-60^{\circ}$, $\theta_{2}=-30^{\circ}$, $\theta_{3}=30^{\circ}$, and $\theta_{4}=60^{\circ}$, respectively. The RCS of the target is set as $\sigma_{0}^{RCS}=20$ $\mathrm{m^{2}}$. According to [22], the mmWave channel complex gain $\beta_{\ell,k}$ follows a complex Gaussian distribution as follows:
$\beta_{\ell,k}\sim\mathcal{C}\mathcal{N}(0,10^{-0.1\chi}),$ (56)
where $\chi=\chi_{a}+10\chi_{b}\log_{10}({\tilde{d}})+\chi_{c}$. ${\tilde{d}}$ denotes the distance between the BS and the user, and $\chi_{c}\sim\mathcal{C}\mathcal{N}(0,\sigma_{c}^{2})$. For the LOS component, the values of $\chi_{a}$, $\chi_{b}$, and $\sigma_{c}$ are set as $\chi_{a}=61.4$, $\chi_{b}=2$, and $\sigma_{c}=5.8\ \mathrm{dB}$, while for the NLOS component, the values are set as $\chi_{a}=72.0$, $\chi_{b}=2.92$, and $\sigma_{c}=8.7\ \mathrm{dB}$, respectively. Unless mentioned otherwise, we consider the user noise $\sigma_{{k}}^{2}=-100\ \mathrm{dBm},k\in\mathcal{K}$ and the sensing echo noise $\sigma_{{z}}^{2}=-100\ \mathrm{dBm}$. The weight factor is set as $\alpha_{k}=1,k\in\mathcal{K}$, and the input SCNR threshold for target sensing is set as $\overline{\gamma}_{r}^{in}=10\ \mathrm{dB}$. The total budget power of the system is $P_{T}=30\ \mathrm{dBm}$. For the proposed WMMSE-PDD algorithm, we consider the initial penalty parameter $\rho^{(0)}=0.01$ and the control constant $c_{\rho}=0.8$. In addition, we consider $\zeta^{(0)}=0.01$ and, in the practical simulation, set $N^{in}_{max}=50$ to ensure the convergence of the inner loop. The accuracy tolerance $\epsilon^{out}$ is set to $10^{-6}$, and $h(\mathbb{X}^{(m)})\leq\epsilon^{out}$ is the condition for the end of Algorithm 2.
Further, to compare the performance of the system under different mmWave channel conditions, we define two types of channels as follows:
(1) Low spatial correlation channel: the LoS components of user channels are located at different spatial angles;
(2) High spatial correlation channel: the LoS components of user channels are located at the same spatial angle.
In the simulation, for type (1), the spatial angles of the LOS components for the four users are set to be $\theta_{1,1}=-40^{\circ}$, $\theta_{1,2}=-15^{\circ}$, $\theta_{1,3}=20^{\circ}$, and $\theta_{1,4}=45^{\circ}$; for type (2), the spatial angles of the LOS components for all users are set to be $\theta_{1,k}=-20^{\circ},k\in\mathcal{K}$. Note that RSMA-assisted mmWave ISAC systems with the above two types of channel scenarios are denoted as “RSMA-ISAC-Low” and “RSMA-ISAC-High,” respectively. All the simulation results are obtained by averaging over 50 channels.
V-A Performance of the proposed WMMSE-PDD Algorithm
Fig. 3 depicts the convergence performance of the proposed WMMSE-PDD algorithm. Specifically, Fig. 3(a) shows the value of the OF (15a) versus the number of outer iterations of Algorithm 2 for two distinct types of channel conditions. It can be seen that both converge within 20 iterations. Fig. 3(b) illustrates the corresponding value of the constraint violation, which can be seen to be below $10^{-6}$ after 60 iterations. It can be seen from the figures that the proposed WMMSE-PDD algorithm has rapid convergence and is feasible for solving the original problem. Furthermore, the significant difference in WSR under two distinct types of channel conditions suggests that LoS components with different spatial angles can provide greater spatial degrees of freedom (DoF) to the user and improve the communication performance of the system. To verify the superiority of the proposed WMMSE-PDD algorithm, we compare it in Fig. 3 with the following benchmark algorithms: 1) the Consensus-ADMM algorithm [30], which is a distributed optimization method; and 2) the “Two-Stage” algorithm[25], which seeks to minimize the Euclidean distance of the optimal fully-digital beamforming matrix. It can be seen from Fig. 3 that, the proposed algorithm outperforms the above two benchmark algorithms in terms of the communication-sensing achievable performance region.
V-B Performance Comparison of Different Schemes
In this subsection, to verify the superiority of the proposed RSMA-assisted mmWave ISAC scheme, we consider the benchmark schemes for comparison as follows:
Figure 2: Convergence performance of the proposed WMMSE-PDD algorithm.
@@ V-B Performance Comparison of Different Schemes
  • 1)
    SDMA-assisted ISAC: This scheme is based on a conventional multiple-access strategy with multi-user linear precoding (MU-LP) [34], in which each user decodes the desired message by treating the interference from the other user’s data streams as noise. Specifically, the SDMA-assisted ISAC scheme can be realized by disabling the common data stream in Eq. (1).
  • 2)
    NOMA-assisted ISAC: This scheme is based on a multiple access strategy with power domain superposition coding at the transmitter side and SIC at the user side [46]. Specifically, the decoding order $\pi$ is obtained based on the descending order of user channel strengths. Thus, the message of user-$\pi(k)$, $k\geq j$, is decoded at user-$\pi(j)$ using the SIC, and the corresponding SINR, denoted by $\gamma_{\pi(j)\rightarrow\pi(k)}$, can be given by
    $\gamma_{\pi(j)\rightarrow\pi(k)}=\frac{\left|\mathbf{h}_{\pi(j)}^{H}\mathbf{F}_{\rm RF}\mathbf{f}_{{\rm D},\pi(k)}\right|^{2}}{\sum\limits_{i<k,i\in\mathcal{K}}\left|\mathbf{h}_{\pi(j)}^{H}\mathbf{F}_{\rm RF}\mathbf{f}_{{\rm D},\pi(i)}\right|^{2}+\sigma_{{j}}^{2}}.$ (57)
    Further, the HBF problem for NOMA-assisted ISAC can be given as follows:
[…]
    where $R_{\pi(k)}=\min_{j\leq k,j\in\mathcal{K}}\{\log_{2}(1+\gamma_{\pi(j)\rightarrow\pi(k)})\}$. Problem P4 is still solved using the proposed WMMSE-PDD algorithm.
@@
Fig. 4 shows user WSR versus budget power for different schemes with fixed sensing SCNR. For low spatial correlation channels, it can be seen that the WSR of the RSMA-ISAC scheme is slightly higher than that of the SDMA-ISAC scheme but significantly higher than that of the NOMA-ISAC scheme. Its reason is that under the high spatial DoF of mmWave channels configured with massive antennas, the strategy that makes sure that the strong user can fully decode the weak user message in the NOMA scheme will lead to the loss of multiplexing gain and rate. For the high spatial correlation channel, it can be seen that at lower budget power, the WSR of the schemes is ranked from high to low as follows: RSMA-ISAC, NOMA-ISAC, and SDMA-ISAC, but as the budget power increases, the SDMA-ISAC scheme outperforms the NOMA-ISAC. In summary, the common data stream of our proposed RSMA-ISAC scheme can reduce interference between users and further improve WSR.
Figure 6: Transmit beampatterns for low spatial correlation channels.
Figure 7: Transmit beampatterns for high spatial correlation channels.
Fig. 5 depicts user WSR versus sensing SCNR thresholds for different schemes. First, when the sensing SCNR threshold is lower than $0\ {\rm dB}$, the WSR remains constant for all schemes, which indicates that there are sufficient resources to serve multiuser communication with low sensing performance requirements. Next, for the two distinct types of channel scenarios, it is observed that RSMA-ISAC schemes both achieve the largest communication-sensing performance region, which indicates that the RSMA scheme’s common data stream can also further manage the interference between radar sensing and communication, providing a greater DoF for the system’s sensing function. Note that when the sensing performance is extremely demanding, i.e., $\gamma_{0}=15\ {\rm dB}$ in the figure, the WSR of the NOMA-ISAC scheme outperforms both RSMA-ISAC and NOMA-ISAC, which is attributed to the fact that the SIC decoding strategy is used for each user in the NOMA scheme, which also provides a certain performance gain in the case of scarce communication resources. Overall, our proposed RSMA-ISAC scheme achieves better performance trade-offs and the largest achievable performance region.
Figs. 6 and 7 show the transmit beampatterns of the different schemes under the power budget $P_{T}=30\ {\rm dBm}$ and the sensing SCNR threshold $\gamma_{0}=10\ {\rm dB}$. Specifically, the transmit beampattern is defined as follows: $P(\theta)=10\log_{10}\big{(}\mathbf{a}_{t}^{H}(\theta)\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm F}_{\rm D}\boldsymbol{\rm F}_{\rm D}^{H}\boldsymbol{\rm F}_{\rm RF}^{H}\mathbf{a}_{t}(\theta)\big{)}$. For the low spatial correlation channel, Fig. 6 depicts that all three different schemes are able to form high-gain beams on the target spatial angle $\theta_{0}=10^{\circ}$, but only the RSMA-ISAC scheme has the deepest beam depression in the four clutter directions. In addition, due to the sparsity of the mmWave channel, directional beams are also formed in the spatial directions of the four user LoS components, but among them, the user beams of the RSMA-ISAC are more stable than those of the NOMA. For the high spatial correlation channel, Fig. 7 similarly depicts that high gain beams can be formed in the target direction, but the SDMA-ISAC scheme has the weakest resistance to clutter interference. And a directional beam is formed in the direction where the spatial angles of all four user LoS components are the same, with RSMA-ISAC having the best beam performance. In summary, the RSMA-ISAC scheme has stronger anti-clutter interference and more stable multi-beam formation capability.
Figure 8: Normalized beampattern of the common data streams.
Fig. 8 depicts the spatial power distribution of the common data stream in the RSMA-ISAC scheme to further reveal the role of the common data stream in the ISAC system. The normalized beampattern is defined as follows:
$P(\theta_{i})=\frac{\left|\mathbf{a}_{t}^{H}(\theta_{i})\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm f}_{\rm c}\right|^{2}}{\max_{\theta_{i}\in\Theta}\left|\mathbf{a}_{t}^{H}(\theta_{i})\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm f}_{\rm c}\right|^{2}},\theta_{i}\in\Theta,$ (59)
where $\Theta=[-\frac{\pi}{2}:\frac{\pi}{180}:\frac{\pi}{2}]$ is the angle grid. For low spatial correlation channels, the common data stream concentrates on the power mainly in the direction of the target spatial angle, which further improves the sensing capability of the ISAC system. On the other hand, for high spatial correlation channels, the common data stream concentrates on the power in the spatial angle where all the user LoS components are the same, i.e., $\theta_{1,k}=-20^{\circ},k\in\mathcal{K}$, which serves to mitigate the interference among the dense users. Therefore, the above analysis further verifies that the common data stream has the role of both mitigating multiuser interference and the trade-off between communication and sensing performance.
Figure 9: WSR $R_{\rm WSR}$ versus sensing SCNR $\gamma_{0}$.
Figure 10: Energy efficiency versus sensing SCNR $\gamma_{0}$.
V-C Hybrid Beamforming Performance of RSMA-ISAC with Finite-Resolution Phase Shifters
In this subsection, we analyse the trade-offs between the WSR and energy efficiency of the communication versus the sensing SCNR thresholds in the RSMA-ISAC system under different numbers of RF chains and phase shifter phase quantization bits, respectively. Energy efficiency is defined as the ratio between WSR and total power consumption and is given by [47]:
${\rm e}=\frac{R_{\rm WSR}}{P_{T}+N_{\rm RF}P_{\rm RF}+N_{\rm PS}P_{\rm PS}(b)+P_{\rm BB}}({\rm{bps/Hz/W}}),$ (60)
where $P_{\rm RF}$ denotes the power consumption of each RF chain, $P_{\rm PS}(b)$ is the power consumption of each PS with $b$-bit resolution, and $P_{\rm BB}$ is the baseband power consumption. $N_{\rm PS}$ is the total number of PSs, where $N_{\rm PS}=0$ for fully digital structures and $N_{\rm PS}=N_{t}N_{\rm RF}$ for hybrid structures. In the simulation, we adopt typical values of $P_{\rm RF}=300\ {\rm mW}$ and $P_{\rm BB}=200\ {\rm mW}$ [47]. The power consumption values of each PS are $2.5$, $5$, and $11.7\ {\rm mW}$ for 1-, 2-, and 4-bit resolution phase shifting [48]. Then $P_{\rm PS}(\infty)=17.6\ {\rm mW}$ for each infinite-resolution PS [49].
From Fig. 9, it can be seen that the WSR versus SCNR achievable performance region of the RSMA-ISAC system based on the WMMSE-PDD algorithm becomes larger with the increase in the number of RF chains and the number of phase quantization bits. When $b$ tends to infinity, the increase of RF chains improves WSR very little, but the achievable performance region of the hybrid structure has reached more than $98\%$ of the fully digital structure. Next, From Fig. 10, it is observed that the energy efficiencies at fixed perceived SCNR thresholds are consistently higher for the scenario with the number of RF chains $N_{\rm RF}=6$ compared to $N_{\rm RF}=10$, and all hybrid structures have higher energy efficiency than the fully digital structure. In this example, when the sensing SCNR is less than $11\ {\rm dB}$, the energy efficiency of the communication increases with the decrease in the number of phase quantization bits, but when the SCNR is greater than $13\ {\rm dB}$, the energy efficiency of the case “$N_{\rm RF}=6,b=\infty$” is the highest. In summary, the RSMA-ISAC with the HBF structure can achieve favorable performance trade-offs and greatly improve the energy efficiency of the system.
VI Conclusions

### v2

@@ III-D Summary of the Proposed WMMSE-PDD Algorithm
1: Define the accuracy tolerance $\epsilon^{out}$ and the maximum number of iterations $N_{max}^{out}$. Initialize $\mathbb{X}^{(0)}$, $\boldsymbol{\rm\Lambda}^{(0)}$, $\rho^{(0)}$, $\zeta^{0}$, set $0<c_{\rho}<1$, the iteraion number $m=0$.
@@
3:  Update $\mathbb{X}^{(m+1)}$ according to Section III-C;
4:  if $h(\mathbb{X}^{(m+1)})\leq\zeta^{(m)}$ then
@@
IV Numerical Results
In this section, we provide numerical results to demonstrate the performance of the proposed RSMA-enhanced mmWave ISAC system with HBF structure. Throughout the whole simulation, we set the carrier frequency as 28 $\mathrm{GHz}$, $N_{t}=32$, $N_{r}=32$, and $K=4$. Unless mentioned otherwise, we consider the user noise $\sigma_{{k}}^{2}=-100\ \mathrm{dBm}$ and the sensing echo noise $\sigma_{{z}}^{2}=-100\ \mathrm{dBm}$. The weight factor is set as $\alpha_{k}=1$, and the input SCNR threshold for target sensing is set as $\overline{\gamma}_{r}^{in}=10\ \mathrm{dB}$. The system power budget is $P_{T}=30\ \mathrm{dBm}$. Further, to compare the performance of the system under different mmWave channel conditions: low spatial correlation channel and high spatial correlation channel.
@@ IV Numerical Results
To verify the superiority of the proposed WMMSE-PDD algorithm, we compare it in Fig. 2 with the following benchmark algorithms: 1) the Consensus-ADMM algorithm [16][17], which is a distributed optimization method; and 2) the “Two-Stage” algorithm[11], which seeks to minimize the Euclidean distance of the optimal fully-digital beamforming matrix. It can be shown that, the proposed algorithm outperforms the above two benchmark algorithms in terms of the communication-sensing achievable performance region.
@@
Figures 3 and 4 show user WSR versus budget power and sensing SCNR threshold, respectively. For both high and low spatial correlation channels, the RSMA-ISAC scheme outperforms the SDMA-ISAC and NOMA-ISAC schemes in terms of WSR and anti-clutter interference capability. We reveal that the common data stream has a dual function in mitigating multi-user interference and trade-offs between communication and sensing performance. For low spatial correlation channels, it can be seen that the WSR of the RSMA-ISAC scheme is slightly higher than that of the SDMA-ISAC scheme but significantly higher than that of the NOMA-ISAC scheme. The reason is that under the high spatial DoF of mmWave channels configured with massive antennas, the strong user can fully decode the weak user message in the NOMA scheme, thus leading to the loss of multiplexing gain and rate. In summary, the common data stream of our proposed RSMA-ISAC scheme can reduce interference between users and further improve WSR. Note that when the sensing performance is extremely demanding, i.e., $\gamma_{0}=15\ {\rm dB}$ in the Fig.4, the WSR of the NOMA-ISAC scheme outperforms both RSMA-ISAC and NOMA-ISAC, which is attributed to the fact that the SIC decoding strategy is used for each user in the NOMA scheme, which also provides a certain performance gain in the case of scarce communication resources.
V Conclusions